OpenStudy (isabel☺):

Find the volume: y=x^2 + 4, y=2x^2

OpenStudy (anonymous):

On what interval? The volume is infinite otherwise.

OpenStudy (isabel☺):

no given

OpenStudy (anonymous):

The x^2+4 is important. Either you updated it or I didn't see it. It's important because we want to find out where the graphs intersect to know our limits of integration. Start by setting them equal, then solve for x to find the points of intersection. x^2+4=2x^2 Solve this for x and you find that they intersect at x = 2 and x = -2, So we integrate from 2 to -2 Let F(x)=x^2+4 and G(x)=2x^2 In this case F(x) is our outer radius and G(x) is our inner radius. Use the ring method to do this. Evaluate the integral: \[\pi*\int\limits_{-2}^{2}(F(x)^2-G(x)^2)dx\] and that will give you the volume. Look in your textbook for how the washer method for finding the volume of a solid of revolution works. If anything is unclear, ask something specific.

OpenStudy (isabel☺):

does the graph looks like this? i really can't find the volume